OpenStudy (anonymous):

You want to determine the number of text messages students at your school send in a month. You randomly ask students in each of your classes. A.random B.systematic C.stratified D.none of these

OpenStudy (anonymous):

I was just going to tag u

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

It is random it seems like haha :D

OpenStudy (anonymous):

A yogurt shop offers 4 different flavors of frozen yogurt and 8 different toppings. How many choices are possible for a single serving of frozen yogurt with 1 topping? 32 16 64 1,680

OpenStudy (anonymous):

I believe for this one, you just multiply 4*8 resulting in 32

OpenStudy (anonymous):

If you choose a topping, you have 4 different choices for that one topping.

OpenStudy (anonymous):

Since there are 8 toppings and there are 4 choices for each topping. 8*4 = 32

OpenStudy (anonymous):

A drug trial is testing the effectiveness of two drugs. If 40 patients are given Drug A, 80 patients are given Drug B, and 50 patients are given a placebo, what is the probability that a patient will not be given a placebo? (1 point) A:8/17 B:12/17 C:4/17 D:5/17

OpenStudy (anonymous):

How many patients are there in all?

OpenStudy (anonymous):

170

OpenStudy (anonymous):

Ok, and how many patients are given a placebo?

OpenStudy (anonymous):

50

OpenStudy (anonymous):

Right! So 50 of the 170 patients are given a placebo.

OpenStudy (anonymous):

So logically, this means that 120/170 patients are NOT given a placebo.

OpenStudy (anonymous):

This will reduce to 12/17

OpenStudy (anonymous):

ty i have a couple more

OpenStudy (anonymous):

Suppose you flip a coin and roll a number cube. What is the probability of having tails come up on the coin and an even number on the number cube?

OpenStudy (anonymous):

Alright, I'll try. Don't be surprised if I don't know one of them. I haven't done stats in a while.

OpenStudy (anonymous):

A: 1/4 B: 1/2 C: 1 D:3/4

OpenStudy (anonymous):

its fine...

OpenStudy (anonymous):

Ok. So the probability of the coin being tails is \[\frac{ 1 }{ 2 }\] and the probability of the dice rolling an even number is also \[\frac{ 1 }{ 2 }\] All we have to do is multiply both of these probabilities together. \[\frac{ 1 }{ 2 }*\frac{ 1 }{ 2 }\]

OpenStudy (anonymous):

A

OpenStudy (anonymous):

\[\frac{ 1 }{ 2 } *\frac{ 1 }{ 2 } = \frac{ 1 }{ 4 }\]

OpenStudy (anonymous):

B?

OpenStudy (anonymous):

So there are 6 possible places that the spinner can land. Of those 6 possible places, there are 4 dark spots.

OpenStudy (anonymous):

2:4

OpenStudy (anonymous):

It would actually be \[\frac{ 4 }{ 6 } = \frac{ 2 }{ 3 }\] Do you understand why?

OpenStudy (anonymous):

yea cause theres 6 total 4 dark spots and you dived to get that

OpenStudy (anonymous):

ohh it was A!

OpenStudy (anonymous):

It should be 2 to 3

OpenStudy (anonymous):

How did you get A?

OpenStudy (anonymous):

idk it says that was the answer...

OpenStudy (anonymous):

who knows man

OpenStudy (anonymous):

OH!

OpenStudy (anonymous):

There are simply twice as many dark spots as there are white spots. Therefore, 2 to 1 odds. haha.

OpenStudy (anonymous):

Ohh haha .You think you could help with this?

OpenStudy (anonymous):

Or this one?

OpenStudy (anonymous):

For the second page its 1-3 with work...

OpenStudy (anonymous):

For the first one. 1.It wouldn't be representative of your school because you're only in one place. The gym at that. 2. Yes, the location could be somewhere where there is a more diverse group of students. 3. Your survey would obviously change because you are being biased. You are obviously judging the people who don't exercise negatively. Idk, man most of this stuff is subjective. You kind of have to make it up as you go.

OpenStudy (anonymous):

I gtg though man. It's pretty late. cya!

OpenStudy (anonymous):

awwh could u help tmrw?